# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

### v1

@@ Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 53, 81. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 219.
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 128. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 227. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 58. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 53, 81, 227.
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 230. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 131, 184.
Concurrent to our survey, several related surveys focused on biomedical-related LLMs. For example, Liu et al. 135 review medical LLMs and He et al. 81 discuss LLMs for healthcare. Similarly, Xiao 227 conducted a comprehensive overview of LLMs in medicine. Other discussed RAG systems in general domains, such as Fan et al. 53 and Gao et al. 65 focused on the architectures, training strategies, and application of RAG in general domains.
@@
LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 247. Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 197. RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 65. This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 227.
In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources. The re-ranking method then sorts these references based on their relevance. Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response. As Fig. 1 shows, the basic framework of biomedical RAG consists of four key components 53: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
@@
RAG has been applied to a variety of biomedical applications. In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 141, enhances specialized decision-making through expert-guided LLMs 119, and facilitates rare disease diagnosis 240. This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 64 and advances genomic medicine through improved variant annotation and interpretation 142. In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 177. Across these applications, RAG leverages diverse data sources—including health records 105, 83, scientific literature 37, clinical guidelines 21, and genomic databases 142—to generate evidence-based responses that significantly improve both clinical care and research outcomes. The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
@@
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 99, 243 and dense retrieval methods 168, 76, LLM-based retrieval 128, 163, knowledge graph enhancement techniques 108, 237, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 199, 177, 215. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 97, 94, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 165, 18, 73 and open-source models 7, 204, 156, 44 alongside domain adaptation techniques 58, 19 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
@@
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 65. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 218, 203. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains190. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this section, we describe various retrieval methodologies. We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
@@
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) 182 and TF-IDF (Term Frequency-Inverse Document Frequency) 106, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 42, 71 implementations to hybrid approaches 120, 13, 3, 115 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 109. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 234, 88, 92. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch 51, Whoosh 206, and Entrez API 56 also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 231, 120, 13, 3, 115, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
@@
Clinfo.ai 141 Entrez API Medical QA 2023
MEDRAG toolkit 231 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS 88 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System 120 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA 13 BM25 and MedCPT Medical Question Answering 2024
CliniqIR 3 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery 115 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT 142 Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT 71 BM25 algorithm Medical QA 2025
Two-Layer RAG 42 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt 169 Entrez API for PubMed retrieval Scientific data analysis 2025
@@
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 246. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 190. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 183, and Sentence Transformer 207 have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding 228, Dense Passage Retriever 111 and GIST Large Embedding 193 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText 26, RoBERTa 137, and SimCSE 63, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
@@
AskFDALabel 19 Sentence Transformer 207 FDA drug labeling extraction 2024
CaLM 168 BGE-embedding 228 with Chroma DB Supporting caregivers FM 2024
RAG-HPO 67 FastEmbed 205 Rare genetic disorders automated deep phenotyping 2024
RALL 76 Dense Passage Retriever 111 Lay language generation 2024
RT 129 BERT embeddings 46 Few-shot medical NER 2024
RAG-GPT 233 BGE-embedding 228 Breast cancer nursing care QA 2024
GNQA 41 HNSW graphs 147 Medical QA 2024
Klang et al. 116 GIST Large Embedding 193 with FAISS ICD-10-CM coding 2024
RAGPR 248 SBERT embeddings 183 Personalized physician recommendations 2025
JGCLLM 61 GLuCoSE-base-ja vectors 60 Genetic counseling support 2025
@@
DRAGON-AI 209 OpenAI Text-Embedding 164 with Chroma DB Ontology generation 2024
RISE 217 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG 121 UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT 130 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat 59 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT 139 Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT 198 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER 214 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP 20 Amazon Titan Text Embeddings v2 14 Nutrition-related question answering 2025
InfectA-Chat 186 Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG 107 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
@@
PM-Search 99 BioBERT 123 Clinical literature retrieval 2022
LADER 98 PubMedBERT 75 Biomedical literature retrieval 2023
CLEAR 140 BioBERT Clinical information extraction 2025
KG-RAG 194 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule 6 MedCPT, BioBERT, BioGPT 145 with Pinecone DB Biomedical QA 2024
@@
BiomedRAG 128 MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM 163 GPT4All 16 with Chroma DB Surgical decision support 2024
Myers et al. 160 LLM2Vec (for comparison) 22 Clinical Information retrieval 2024
@@
Self-BioRAG 93 MedCPT 97 Medical QA 2024
WeiseEule 6 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR 3 MedCPT and BM25 Diagnostic decision support 2024
MEDRAG 6 BM25, SPECTER 37, Contriever 92and MedCPT Biomedical RAG Tool 2024
@@
RAMIE 242 MedCPT, Contriever and BMRetriever 234 Biomedical IR about dietary supplements 2025
@@
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings dominate, alternatives are emerging. The RAP system 20 utilizes Amazon’s Titan Text Embeddings v2 14 for nutrition-related question answering. Other commercial APIs such as Anthropic’s Voyage-2 17 and Google’s Vertex AI embeddings 74, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
@@
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 123, 75. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 146. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 146. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 75, BioBERT 123 and BioGPT 145, which have been widely adopted in biomedical RAG systems. For example, PM-Search 99 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER 98 for biomedical literature retrieval and in KG-RAG 194 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule 6 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 97 to achieve enhanced performance in biomedical QA.

### v2

@@ Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 1, 2. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 3.
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 4. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 5. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 6. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 1, 2, 5.
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 7. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 8, 9.
Concurrent to our survey, several related surveys focused on biomedical-related LLMs. For example, Liu et al. 10 review medical LLMs and He et al. 2 discuss LLMs for healthcare. Similarly, Xiao 5 conducted a comprehensive overview of LLMs in medicine. Other discussed RAG systems in general domains, such as Fan et al. 1 and Gao et al. 11 focused on the architectures, training strategies, and application of RAG in general domains.
@@
LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 12. Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 13. RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 11. This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 5.
In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources. The re-ranking method then sorts these references based on their relevance. Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response. As Fig. 1 shows, the basic framework of biomedical RAG consists of four key components 1: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
@@
RAG has been applied to a variety of biomedical applications. In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 14, enhances specialized decision-making through expert-guided LLMs 15, and facilitates rare disease diagnosis 16. This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 17 and advances genomic medicine through improved variant annotation and interpretation 18. In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 19. Across these applications, RAG leverages diverse data sources—including health records 20, 21, scientific literature 22, clinical guidelines 23, and genomic databases 18—to generate evidence-based responses that significantly improve both clinical care and research outcomes. The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
@@
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 24, 25 and dense retrieval methods 26, 27, LLM-based retrieval 4, 28, knowledge graph enhancement techniques 29, 30, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 31, 19, 32. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 33, 34, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 35, 36, 37 and open-source models 38, 39, 40, 41 alongside domain adaptation techniques 6, 42 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
@@
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 11. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 43, 44. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains45. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this section, we describe various retrieval methodologies. We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
@@
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) 46 and TF-IDF (Term Frequency-Inverse Document Frequency) 47, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 48, 49 implementations to hybrid approaches 50, 51, 52, 53 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 54. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 55, 56, 57. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch 58, Whoosh 59, and Entrez API 60 also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 61, 50, 51, 52, 53, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
@@
Clinfo.ai 14 Entrez API Medical QA 2023
MEDRAG toolkit 61 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS 56 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System 50 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA 51 BM25 and MedCPT Medical Question Answering 2024
CliniqIR 52 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery 53 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT 18 Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT 49 BM25 algorithm Medical QA 2025
Two-Layer RAG 48 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt 62 Entrez API for PubMed retrieval Scientific data analysis 2025
@@
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 63. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 45. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 64, and Sentence Transformer 65 have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding 66, Dense Passage Retriever 67 and GIST Large Embedding 68 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText 69, RoBERTa 70, and SimCSE 71, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
@@
AskFDALabel 42 Sentence Transformer 65 FDA drug labeling extraction 2024
CaLM 26 BGE-embedding 66 with Chroma DB Supporting caregivers FM 2024
RAG-HPO 72 FastEmbed 73 Rare genetic disorders automated deep phenotyping 2024
RALL 27 Dense Passage Retriever 67 Lay language generation 2024
RT 74 BERT embeddings 75 Few-shot medical NER 2024
RAG-GPT 76 BGE-embedding 66 Breast cancer nursing care QA 2024
GNQA 77 HNSW graphs 78 Medical QA 2024
Klang et al. 79 GIST Large Embedding 68 with FAISS ICD-10-CM coding 2024
RAGPR 80 SBERT embeddings 64 Personalized physician recommendations 2025
JGCLLM 81 GLuCoSE-base-ja vectors 82 Genetic counseling support 2025
@@
DRAGON-AI 83 OpenAI Text-Embedding 84 with Chroma DB Ontology generation 2024
RISE 85 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG 86 UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT 87 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat 88 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT 89 Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT 90 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER 91 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP 92 Amazon Titan Text Embeddings v2 93 Nutrition-related question answering 2025
InfectA-Chat 94 Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG 95 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
@@
PM-Search 24 BioBERT 96 Clinical literature retrieval 2022
LADER 97 PubMedBERT 98 Biomedical literature retrieval 2023
CLEAR 99 BioBERT Clinical information extraction 2025
KG-RAG 100 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule 101 MedCPT, BioBERT, BioGPT 102 with Pinecone DB Biomedical QA 2024
@@
BiomedRAG 4 MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM 28 GPT4All 103 with Chroma DB Surgical decision support 2024
Myers et al. 104 LLM2Vec (for comparison) 105 Clinical Information retrieval 2024
@@
Self-BioRAG 106 MedCPT 33 Medical QA 2024
WeiseEule 101 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR 52 MedCPT and BM25 Diagnostic decision support 2024
MEDRAG 101 BM25, SPECTER 22, Contriever 57and MedCPT Biomedical RAG Tool 2024
@@
RAMIE 107 MedCPT, Contriever and BMRetriever 55 Biomedical IR about dietary supplements 2025
@@
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings dominate, alternatives are emerging. The RAP system 92 utilizes Amazon’s Titan Text Embeddings v2 93 for nutrition-related question answering. Other commercial APIs such as Anthropic’s Voyage-2 108 and Google’s Vertex AI embeddings 109, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
@@
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 96, 98. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 110. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 110. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 98, BioBERT 96 and BioGPT 102, which have been widely adopted in biomedical RAG systems. For example, PM-Search 24 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER 97 for biomedical literature retrieval and in KG-RAG 100 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule 101 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 33 to achieve enhanced performance in biomedical QA.